A solution is prepared by mixing 2 g of FeCl3 with enough water to make 50 mL of solution. Calculate the molarity of the solution.


1
Expert's answer
2021-05-20T07:51:16-0400

First we use the molar mass (MM) of FeCl3 [MM = (35.45 g/mol)(3) + (55.85 g/mol)(1) = 162.20 g/mol] and the mass of solid (m) given to calculate the number of moles dissolved:


  • "n = \frac{m}{MM} = \frac{2\,\cancel{g}}{162.20\,\cancel{g}/mol} = 0.01233 \,mol"


The molarity of the solution (M) can be calculated with the number of moles of solute (n) and the volume of the solution (V) already given (remember that the volume has to be expressed in liters):


  • "M = \frac{n}{V} = \frac{0.01233\,mol}{50\, mL} = \frac{0.01233\,mol}{0.05\, L} = 0.2466 \,mol/L"

To conclude, it is found that the molarity of FeCl3 on the solution is 0.247 M or 0.247 mol/L